OpenStudy (anonymous):

Please help i need to see this worked out. It confuses me. 1/x+1/a=1/b a does not =b solve for x

OpenStudy (anonymous):

\[x=(ab)/(a+b)\]

OpenStudy (anonymous):

how did you get the answer

OpenStudy (anonymous):

Here's how: 1/x=1/b -1/a 1/x=a/ab-b/ab (common denominator) 1/x=(a-b)/ab (combine terms) x=1b/(a-b) (take reciprocals 0f both sides)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The last line should be \[ab/(a-b)\]